The management of a restaurant has been studying whether or not new customers return



within a month. The collected data reveal that 60% of the new customers have returned.



If 90 new customers dine at the restaurant this month, what is the probability that at least



60 will return next month? Use Normal approximation to Binomial distribution.

1
Expert's answer
2022-04-11T16:25:23-0400

First, we must verify that the following criteria are met:

"np\ge5,\\
n(1-p) \ge5."

In this case, we have:

"np=90\cdot0.6=54,\\
n(1-p)=90\cdot(1-0.6)=36."

Both numbers are greater than 5, so we’re safe to use the normal approximation.


We must apply a continuity correction when calculating probabilities. We want to find "P(X\ge60)". To use the normal distribution to approximate the binomial distribution, we would instead find "P(X\ge59. 5) ."


The mean and standard deviation of the binomial distribution:

"\mu=np=90\cdot0.6=54,\\
\sigma=\sqrt{np(1-p)}=\\
=\sqrt{90\cdot0.6\cdot(1-0.6)}=4.65."


The z-score:

"z=\cfrac{x- \mu} {\sigma}=\cfrac{59. 5- 54} {4.65}=1.18."


"P(X>59.5)=P(Z>1.18)=\\
=1-P(Z<1.18)=1-0.8810=0.1190."